# On Emergent Social World Models — Evidence for Functional Integration of Theory of Mind and Pragmatic Reasoning in Language Models

### v1

@@ A.2.3 Other Datasets
For comparison, we use items from BLiMP (Warstadt et al., 2020), a widely used benchmark on grammatical phenomena in English. BLiMP consists of minimal pairs, where in each pair, one sentence is grammatical and one is ungrammatical. LMs are evaluated by comparing the probabilities they assign to the grammatical and ungrammatical sentences. We use a subset of 10 sample items per category.
@@ A.4 Models
A.5 Additional Results
We show the models’ average accuracies on the syntax benchmark BLiMP, next to the accuracies on ToM and pragmatics in Figure 4. Consistent with the literature, accuracy on the complex domains ToM and pragmatics increases over model size. However, just model size or model type (base or fine-tuned) are not sufficient for explaining away the connection between pragmatic and ToM performance of LMs, as shown via model comparison in Section 2.3.
Figure 4: Average accuracy (y-axis) for the domains pragmatics, theory of mind and syntax (color) by model size (x-axis).
Figure 5: Detailed overview of the accuracy (y-axis) results of all models (x-axis) on the different pragmatic datasets (facets).
Figure 6: Detailed overview of the accuracy (y-axis) results of all models (x-axis) on the different ToM datasets (facets).
Figure 7: Statistical comparison of models with different combinations of up to seven ATOMS predictors with LOO cross-validation, showing the estimated expected log-likelihoods for LOO-CV (y-axis) for different models (x-axis). Higher values indicate a model’s better predictive fit to the behavioral accuracy on ToM datasets. All models contain baseline predictors model family, size, type and dataset type (summarized as “base”). The ATOMS predictors are abbreviated (“bel.” = belief, “des.” = desire, “emo.” = emotions, “int.” = intentions, “know.” = knowledge, “per.” = percepts, “nlc.” = non-literal communication). The best model contains the predictors “base + int. + des. + emo. + per. + nlc”, the worst model is the “base” only model.
@@ Appendix B Prompts for Stimulus Generation
[examples]
The examples again where five randomly selected story items from the original dataset, separated by "—-" and each prefixed by "Story:"
@@ C.1 Details on each Localizer Suite
Table 3: Examples of stimuli for all localizer suites. Dots indicate answer options.
@@ Appendix D Subnetwork Localization Procedure Details
In the ablation experiments (Section 3.5), each model was evaluated on BLiMP (Warstadt et al., 2020), and on the following selected ToM and pragmatics datasets on which the intact models showed an above-chance performance (described in detail above in Section A.2):
@@
GB 0.59 0.49 0.58 0.5 0.37 0.21
(simple) [0.46, 0.72] [0.35, 0.62] [0.46, 0.7] [0.38, 0.63] [-0.06, 0.81] [-0.25, 0.66]
MI 0.18 -0.11 0.16 -0.08 0.12, 0.05
(simple) [0.05, 0.32] [-0.24, 0.02] [0.04, 0.29] [-0.2, 0.05] [-0.31, 0.55] [-0.4, 0.49]
LB + CI 0.14 -0.05 0.21 -0.1 0.06 0.15
(conj.) [0, 0.27] [-0.17, 0.08] [0.09, 0.34] [-0.22, 0.03] [-0.38, 0.5] [-0.3, 0.6]
All 0.22 -0.05 0.29 0.05 0.11 0.18
(simple) [0.09, 0.36] [-0.18, 0.07] [0.17, 0.42] [-0.07, 0.18] [-0.32, 0.56] [-0.28, 0.63]
CI 0.08 -0.11 0.16 -0.13 0.09 0.08
(conj.) [-0.05, 0.22] [-0.25, 0.02] [0.04, 0.29] [-0.25, -0.01] [-0.35, 0.53] [-0.39, 0.52]
CI 0.14 -0.05 0.23 -0.05 0.11 0.12
(simple) [0, 0.28] [-0.18, 0.08] [0.11, 0.36] [-0.18, 0.07] [-0.32, 0.54] [-0.34, 0.57]
LB 0.26 0.14 0.36 0.15 0.13 0.22
(conj.) [0.13, 0.39] [0.01, 0.27] [0.23, 0.48] [0.03, 0.27] [-0.3, 0.58] [-0.24, 0.68]
LB 0.34 0.27 0.4 0.22 0.19 0.2
(simple) [0.21, 0.47] [0.15, 0.4] [0.27, 0.52] [0.09, 0.34] [-0.24, 0.62] [-0.24, 0.66]

### v2

@@ A.2.3 Other Datasets
For comparison, we use items from two datasets testing LMs’ general linguistic performance. First, across behavioral evaluations and functional localization, we use BLiMP (Warstadt et al., 2020), a widely known benchmark on grammatical phenomena in English. BLiMP consists of minimal pairs, where in each pair, one sentence is grammatical and one is ungrammatical. LMs are evaluated by comparing the probabilities they assign to the grammatical and ungrammatical sentences. We use a subset of 10 sample items per category. Second, for behavioral evaluations we also use SNLI (Bowman et al., 2015), a natural language inference benchmark in English. It consists of pairs of sentences, a premise and a hypothesis. The task is to identify whether the relation between these two sentences is entailment, contradiction or in neutral relation. The models are evaluated by comparing the conditional probabilities the LM assigns to the labels (e.g., “The hypothesis follows from the premise.”). The label to which the model assigns the highest probability given the sentence pair is the model’s prediction. We use a subset of 670 items, to match the size of the BLiMP dataset.
@@ A.4 Models
A.5 Statistical Analyses
Below, we provide details on how the predictions P1–P3 in Section 2 are operationalized and tested statistically. For P2, we used a Bayesian Beta regression model, which predicts the average accuracy of an LM on a given dataset (dependent variable, bound between 0 and 1) according to a Beta distribution, where the predictor for the mean of the distribution is parameterized by several independent variables. For each of the independent variables (i.e., intercept, model family, model size, model training type, dataset type, dataset domain) respective regression coefficients are estimated. For P2, we hypothesized the following: if computational mechanisms are shared between pragmatics and ToM, there will be no credible differences in the accuracy on ToM vs. pragmatics tasks, when controlling for other factors which may lead to differences independently of these mechanisms (e.g., model size), which will be borne out statistically as a non-credible regression coefficient for the predictor “domain” (ToM or pragmatics), which we report in Section 2.3.
For P3, we used two versions of a Bayesian Beta regression model predicting the independent variable “accuracy on pragmatics“ (for each dataset for each model) with the dependent variables model size, model type, and either the average accuracy on linguistic benchmarks BLiMP and SNLI (model M0), or on ToM (model M1). As standard in leave-on-out cross-validation, we calculated the expected log predictive density (ELPD) of each of the two models, by iteratively fitting each model on the empirical data but a left out data point (done through the R package brms, Bürkner, 2017). We then compared the two models by evaluating the log predictive density of the held-out observation under each model, computing a $p$-value for this using Lambert’s $z$-score method.
A.6 Additional Results
We show the models’ average accuracies on the linguistic benchmarks BLiMP and SNLI, next to the accuracies on ToM and pragmatics in Figure 4. Consistent with the literature, accuracy on the complex domains ToM and pragmatics increases over model size. However, just model size or model type (base or fine-tuned) are not sufficient for explaining away the connection between pragmatic and ToM performance of LMs, as shown via model comparison in Section 2.3.
Figure 4: Average accuracy (y-axis) for the domains pragmatics, theory of mind and general linguistic abilities (color) by model size (x-axis).
Figure 5: Detailed overview of the accuracy (y-axis) results of all models (x-axis) on the first eight pragmatic datasets (facets).
Figure 6: Detailed overview of the accuracy (y-axis) results of all models (x-axis) on the remaining eight pragmatic datasets (facets).
Figure 7: Detailed overview of the accuracy (y-axis) results of all models (x-axis) on ten of the ToM datasets (facets).
Figure 8: Detailed overview of the accuracy (y-axis) results of all models (x-axis) on the remaining twelve ToM datasets (facets).
Figure 9: Statistical comparison of models with different combinations of up to seven ATOMS predictors with LOO cross-validation, showing the estimated expected log-likelihoods for LOO-CV (x-axis) for different regression models (y-axis). Higher values indicate a model’s better predictive fit to the behavioral accuracy on ToM datasets. All models contain baseline predictors model family, size, type and dataset type (summarized as “base”). The ATOMS predictors are abbreviated (“bel.” = belief, “des.” = desire, “emo.” = emotions, “int.” = intentions, “know.” = knowledge, “per.” = percepts, “nlc.” = non-literal communication). The best model contains the predictors “base + int. + des. + emo. + per. + nlc”, the worst model is the “base” only model.
@@ Appendix B Prompts for Stimulus Generation
{examples}
The examples again were five randomly selected story items from the original dataset, separated by "—-" and each prefixed by "Story:"
@@ C.1 Details on each Localizer Suite
Table 3: Examples of synthetic GPT-5 generated stimuli for all localizer suites. Dots indicate answer options. The references indicate the source of the original stimuli that were used to prompt GPT-5.
@@ C.2 Formatting in Localization Procedures
C.3 Principal Component Analyses of Synthetic Localizers
To assess the GPT-5 generated localizer stimuli in comparison to original localizer materials from the literature in terms of surface-level and statistical properties to which LMs might be particularly sensitive, a Principal Component Analysis (PCA) with two components was performed. For each condition of each suite, sentence embeddings for the story part (i.e., the context narrative without the question or answer options) of all original stimuli and an equal number of randomly selected synthetic stimuli were calculated using the SentenceTransformer’s all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019). The results of the PCA analysis of the original and synthetic data for each condition are showing in Figures 1011. Visual inspection confirms that original and synthetic stimuli do not cluster separately, indicating that there is no systematic mismatch between them.
Only for one of the conditions (HumanDescr from the LatentBeliefs suite), the two groups can clearly be separated; manual inspection suggests that the difference lies in the fact that the original stimuli describe the physical appearance of a person only, while the synthetic stimuli include other characteristics such as habits. Since the task at hand (i.e., answering simple questions about the described person) does not require Theory-of-Mind reasoning in either case, this does not affect using the synthetic stimuli as a non-ToM control condition. For two other conditions (FalsePhotograph from the LatentBeliefs suite and DecOutcome from the MoralIntent suite), some weak clustering is visible, but original and synthetic stimuli are not clearly separated, and the amount of clustering differs strongly between different samples of synthetic stimuli. Finally, for GameBelief and GameOutcome from the GameBeliefs suite, the sentence embeddings form two clusters, both including a similar number of original and synthetic stimuli; this can be explained by the presence of two types of strategic games in the dataset (trust games and ultimatum games).
Figure 10: PCA with two components of sentence embeddings (from SentenceTransformer’s all-MiniLM-L6-v2 model) of the original stimuli and an equal number of randomly selected synthetic stimuli, for the conditions FalseBelief, Desire, FalsePhotograph, HumanDescr, MechInf, and NonhumanDescr (all from the LatentBeliefs suite).
Figure 11: PCA with two components of sentence embeddings (from SentenceTransformer’s all-MiniLM-L6-v2 model) of the original stimuli and an equal number of randomly selected synthetic stimuli, for the conditions Deceptive, Ironic, Literal, Meaningless (from the CommunicativeIntent suite), GameBelief, GameOutcome (from GameBeliefs), MoralIntent, DecOutcome (from MoralIntent, respectively).
@@ Appendix D Subnetwork Localization Procedure Details
In the ablation experiments (Section 3.5), each model was evaluated on BLiMP (Warstadt et al., 2020) and SNLI (Bowman et al., 2015), and on the following selected ToM and pragmatics datasets on which the intact models showed an above-chance performance (described in detail above in Section A.2):
@@
GB 0.59 0.49 0.58 0.50 0.36 0.22
(simple) [0.46, 0.73] [0.36, 0.61] [0.46, 0.7] [0.39, 0.63] [0.06, 0.64] [-0.09, 0.55]
MI 0.19 -0.11 0.16 -0.08 0.08 0.08
(simple) [0.05, 0.32] [-0.24, 0.02] [0.04, 0.30] [-0.20, 0.05] [-0.22, 0.38] [-0.24, 0.41]
LB + CI 0.14 -0.04 0.21 -0.10 0.04 0.17
(conj.) [0.01, 0.28] [-0.17, 0.09] [0.08, 0.34] [-0.22, 0.03] [-0.26, 0.33] [-0.14, 0.50]
All 0.22 -0.05 0.29 0.05 0.10 0.19
(simple) [0.09, 0.36] [-0.19, 0.08] [0.17, 0.42] [-0.07, 0.17] [-0.19, 0.40] [-0.14, 0.50]
CI 0.08 -0.12 0.16 -0.13 0.03 0.13
(conj.) [-0.05, 0.22] [-0.24, 0.02] [0.04, 0.29] [-0.26, -0.01] [-0.27, 0.33] [-0.19, 0.45]
CI 0.14 -0.05 0.23 -0.05 0.03 0.20
(simple) [0.01, 0.28] [-0.19, 0.08] [0.10, 0.36] [-0.18, 0.07] [-0.26, 0.32] [-0.12, 0.52]
LB 0.26 0.14 0.36 0.15 0.18 0.18
(conj.) [0.13, 0.40] [0.00, 0.27] [0.23, 0.48] [0.02, 0.28] [-0.12, 0.47] [-0.14, 0.50]
LB 0.34 0.27 0.39 0.22 0.20 0.19
(simple) [0.20, 0.47] [0.14, 0.40] [0.26, 0.52] [0.09, 0.34] [-0.10, 0.50] [-0.13, 0.51]
@@
D.1 Subnetwork Ablation Results on Additional Datasets
Next to assessing the effect of the subnetwork ablation on ToM and pragmatics performance in comparison to general linguistic ability benchmarks, we evaluated 13 models on additional cognitive skills that may be recruited in ToM and pragmatic tasks, namely on the entity tracking and analogical reasoning benchmarks.
We cast both benchmarks into multiple choice format. For the entity tracking benchmark from Kim and Schuster (2023), we use the items from the test dataset from their experiment 1, and convert the masked item format into the following sentence completion prompt:
Given the description after “Description:”, write a true statement about all boxes and their contents according to the description after “Statement:”. Format the statement as Box 0 contains the A, Box 1 contains the B, Box 2 contains the C, Box 3 contains the D, Box 4 contains the E, Box 5 contains the F, Box 6 contains the G. A, B, C, D, E, F, G are placeholders for the contents of each box.
Description: Box 0 contains {A}, Box 1 contains {B}, …
Statement: Box
{X} contains
The performance is assessed by retrieving the conditional probability of each of the $m$ different answer options under the LM, where $m$ is the number of boxes described in the item, one of which described the correct content of the target box $X$. We match the size of the linguistic benchmarks and sample 670 items from the original dataset.
For analogical reasoning, the multiple choice formatted verbal analogy and story analogy tasks from Webb et al. (2023) were selected. For verbal analogy, the task is to identify a correct completion which is analogous to an example pair (e.g., given “clothing : jacket :: bird : ”, select among the options pigeon, dog). LMs are evaluated by scoring each of the possible completions, given the example and first member of the target pair. For story analogy, the task is to decide, given a story, which of two additional stories is a better analogy to the first story, or if they are equally analogous. We largely build on the prompts from Webb et al. (2023).
We evaluate the following models on these three datasets: Falcon3-7B, Falcon3-7B-Instruct, Falcon3-10B, Falcon3-10B-Instruct, Llama-3.1-8B, Llama-3.1-8B-Instruct, Llama-2-13b-hf, Llama-2-13b-chat-hf, Llama-3.1-70B, Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-32B, Qwen2.5-32B-Instruct. Each model is evaluated without interventions as well as under ablations of the critical and least active networks localized by each of our localizers. Results for each of the localizers, aggregated across models are shown in Figure 18. For most localizer suites, including the LatentBeliefs (simple & conjunctive) and GameBeliefs suites which mainly supported our overall results reported in Section 3.5, the causal effect of the critical subnetwork ablation was largest for the performance on entity tracking. For analogical reasoning, the effect was smaller (e.g. for GameBeliefs) or qualitatively different (e.g., for CommunicativeIntent conjunctive) than for the critical evaluation domains pragmatics and ToM. This suggests that ToM and pragmatic capabilities might indeed also rely on shared entity tracking mechanisms, in addition to social reasoning mechanisms, and that these may share representations. In contrast, analogical reasoning seems to be less likely to be part of ToM and pragmatics tasks. Results for all 13 models for each localizer are shown in Figures 1920 (aggregating the three reasoning datasets into the “General reasoning” domain).
Figure 18: Average difference in accuracy under ablations of the localized subnetworks, by localizer, across 13 models. Additionally to the results reported in the main text, change in performance on general reasoning tasks (entity tracking Kim and Schuster, 2023; verbal and story analogies Webb et al., 2023) is shown.
Figure 19: Average difference in accuracy under ablations of the localized subnetworks, for each localizer and selected six models. Additionally to the results reported in the main text, average change in performance on control general reasoning tasks is shown.
Figure 20: Average difference in accuracy under ablations of the localized subnetworks, for each localizer and additional selected seven models. Additionally to the results reported in the main text, average change in performance on control general reasoning tasks is shown.